The velocity of a particle at an instant is 10m/s.After 3s it's velocity will become 16 m/s.The velocity at 2s,before the given instant will be

Grade:11

2 Answers

Arun
4 years ago
Acceleration = (v - u) / t
= (16 - 10) / 3
= 2 m/s²
Acceleration of body is 2 m/s²

Now,
v = u + at
10 = u + (2 × 2)
u = 10 – 4
u = 6 m/s
Velocity 3 seconds before was 4 m/s
aswanth nayak
4 years ago
dear student,
 
Acceleration = (v - u) / t= (16 - 10) / 3 = 2 m/s²Acceleration of body is 2 m/s²Now,v = u + at10 = u + (2 × 2)u = 10 – 4u = 6 m/sVelocity 3 seconds before was 4 m/s
Hope it helps you
regards